OpenStudy (anonymous):

Can someone please help with adding and subtracting mixed number im in sixth grade

OpenStudy (anonymous):

what is the problem

OpenStudy (anonymous):

im doin my homwowek i have like 17 left i got stuck on the first

OpenStudy (anonymous):

ok what is the 1st problem

OpenStudy (anonymous):

1.5-3 4/7 and i need shown work

OpenStudy (anonymous):

dude do your own work

OpenStudy (anonymous):

the one isnt part of the prob

OpenStudy (anonymous):

ROBOTDS AND TOAST MAYBE IM STUCK SO SHUT THE HELL UP

OpenStudy (anonymous):

:D is it 1/2 - 3 4/7?

OpenStudy (anonymous):

NOT EVEN CLOSE

OpenStudy (anonymous):

NO IM ASKING IF THATS THE PROBLEM

OpenStudy (anonymous):

or 5 - 3 4/7?

OpenStudy (anonymous):

1.5 converts to 3/2 first

OpenStudy (anonymous):

THERE IS NO 1.5 its just 5

OpenStudy (anonymous):

common denominator, since you have mixed fractions is 7. so 5/1*7/7 = 35/7 then you have 3 4/7. get 3 with a common denominator so 3/1*7/7 = 21/7 then 21/7 + 4/7 = 25/7 then 35/7 - 25/7 = 10/7

OpenStudy (anonymous):

how is it 21/7

OpenStudy (anonymous):

you have 3 and 4/7. so you first have to find a common denominator with 4/7. so you do this by multiplying by 7/7 which is 1 and so 3=3/1*7/7 = 21/7 = 3

OpenStudy (anonymous):

so the answer is ten seveths

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i kneew how to do this but i forgot

OpenStudy (anonymous):

number 2 8-2 and 3/8

OpenStudy (anonymous):

ok so the mixed fraction has a denominator with 8 so use 8 as the common denominator 8*8/8 = 64/8 then the same thing for 2 2*8/8 = 16/8 so 16/8 + 3/8 = 19/8 then 64/8-19/8 = 45/8 and if the answer has to be in mixed fractions the answer is 5 5/8

OpenStudy (anonymous):

nooooooo thts wrong for the second fraction isnt an eight it stays a 3

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

just like i said its 2 and 3/8 not 2 and 8/8

OpenStudy (anonymous):

i know its how i solved it...

OpenStudy (anonymous):

no u SAID 2 AND 8/8

OpenStudy (anonymous):

no 8*8/8 = 64/8 2*8/8 = 16/8 + 3/8 = 19/8 64/8-19/8 = 45/8 = 5 5/8

OpenStudy (anonymous):

next-9 and 3/4-2 and 3/8 and isnt the denominater 8

OpenStudy (anonymous):

yeah try it

OpenStudy (anonymous):

u suck u kno tht if i didnt need help i wouldnt be on here i need help cause i dont get it

OpenStudy (anonymous):

lmao dude you should TRY and then i can help you if you get stuck. that way you LEARN

OpenStudy (anonymous):

i do learn u doggy hiscuit i just did a prob but the denominater was the same

OpenStudy (anonymous):

haha ok ok fine

OpenStudy (anonymous):

so you said common denominator is 8 so then 9*8/8 = 72/8 and 3/4 * 2/2 = 6/8 so 72/8 + 6/8 = 78/8 then 2*8/8 = 16/8 + 3/8 = 19/8 so 78/8 - 19/8 = 59/8 or 7 3/8

OpenStudy (anonymous):

it was 1 and 3/8

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i have a calculator but we have to show our work so i use it to check

OpenStudy (anonymous):

you said 9 3/4 - 2 3/8

OpenStudy (anonymous):

right

OpenStudy (anonymous):

nevermind its seven i put it in wring

OpenStudy (anonymous):

bam!

OpenStudy (anonymous):

SHUT UP AND HELP ME/ NUMBER 6- 6 and 2/3-1 and 1/6

OpenStudy (anonymous):

ok so common denominator is 6 6*6/6 = 36/6 2/3*2/2 = 4/6 so 36/6 + 4/6 = 40/6 then 1*6/6 = 6/6 + 1/6 = 7/6 so 40/6 - 7/6 = 33/6 or 5 1/2

OpenStudy (anonymous):

OK numba 12/ 9 and 6/7-2 and 5/14

OpenStudy (anonymous):

denominaters 14

OpenStudy (anonymous):

ok so denominator is 14 9*14/14 = 126/14 6/7*2/2=12/14 so 126/14+12/14=138/14 then 2*14/14 = 28/14 + 5/14 = 33/14 so 138/14-33/14=105/14 or 7 1/2

OpenStudy (anonymous):

ok number 15. 2 and 2/3-1 and 1/4 denominaters 24

OpenStudy (anonymous):

try it dude

OpenStudy (anonymous):

what did i say about calling me dude u wingspan

OpenStudy (anonymous):

LMAO

OpenStudy (anonymous):

the answer is 1 and 5/12 but i used me calculator hahha feather turned on you forgot about technoglgy

OpenStudy (anonymous):

dude i do this in my head. i dont need calculators

OpenStudy (anonymous):

well im a sixth grader i do get over it

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

done with me math watch out wingspan ill see u in your dreams

OpenStudy (anonymous):

ohhh i gotta make a timeline for reading :( help

OpenStudy (anonymous):

lmao this is math kid! i can help ifyou like though

OpenStudy (anonymous):

dIDNT JADA JUST ASK FOR HELP DIDNT I DIDNT I YES YES YES I DID U SLOW AND RETARDED

OpenStudy (anonymous):

haha whats it on

OpenStudy (anonymous):

have u ever heard of the shut out

OpenStudy (anonymous):

hmmmmmmmmmm no

OpenStudy (anonymous):

well then u cant help me cause its a book about baseball

OpenStudy (anonymous):

dude google

OpenStudy (anonymous):

then your stupid cause u cant read it on google now can u

OpenStudy (anonymous):

yeah i can

OpenStudy (anonymous):

u dotn know what section im on :p